Please answer the question and select upward or downard for each part of the question
Answer #1

(a)

from newtons second law


Fnet = m*a = 92.4*10^-3*15.6 = 1.44 N   upwards

------------------

(b)

Fnet = F - mg


F = Fnet + mg

F = 1.44 + (92.4*10^-3*9.8) = 2.34 N    upwards

----------------

(c)

F = Fg = m*g = 92.4*10^-3*9.8 = 0.905 N downwards

======================


d)


F = Fg = m*g = 92.4*10^-3*9.8 = 0.905 N downwards

acceleration a = F/m = g = 9.81 m/s^2 downwards